### v1

@@ 5 $\Omega(\alpha)$ Informational Gap
In this section we establish a lower bound that exposes a fundamental information gap in the problem of Stochastic Knapsack with Costs. Previous constant-factor approximation algorithms for Stochastic Knapsack (without costs) have used limited distributional information, specifically the expected (or truncated) sizes and the probability of fitting within the current capacity. In Proposition 5.1 we demonstrate that this information is insufficient. We show that any algorithm, even one that finds an adaptive policy, which relies solely on such limited data, cannot achieve a constant-factor approximation in the presence of costs. More generally, we prove in Theorem 5.2 that even providing the first $k$ moments of each distribution for any constant $k$ is insufficient. This showcases the inherent difficulty of the added costs, implying that future constant-approximation algorithms will need to utilize almost the full distributional structure in novel ways.
5.1 Warm-Up: First Moment
Proposition 5.1.
@@ 5.1 Warm-Up: First Moment
Note that in this instance the expected values and the truncated expected values are identical: $\operatorname{\mathbb{E}}[s_{1}]=\operatorname{\mathbb{E}}[s_{2}]=\mu_{1}=\mu_{2}=\epsilon$ and $q_{1}(1)=q_{2}(1)=1$, hence, algorithm $\mathcal{A}$ cannot distinguish between the two items. Inserting one item of type 2 results in a net profit of $\epsilon$, and reduces the remaining capacity $B$ by at least $\epsilon/2$. Inserting any further items (of either types) results in a negative expected profit:
@@
Now, consider an input instance $I$ with $n>\frac{1}{\epsilon}$ items of type 1 and $n^{2}-n$ items of type 2. The optimal policy will attempt to insert items of type 1, and stop if an item’s size instantiates $s_{i}=1$, or if no more items of type 1 remain. The expected number of items successfully inserted by OPT is $\frac{1}{\epsilon}$, each item contributing $\epsilon$ net profit, resulting in
@@
Since $\mathcal{A}$’s only strategy is to randomly select an item to insert, with probability $1-\frac{1}{n}$ it begins by inserting a “bad” item. By taking $n\to\infty$, the expected profit of $\mathcal{A}$ approaches $\epsilon$:
@@ 5.2 $k$ Moments
Inserting a single item of type 2 results in a net profit of $\epsilon$, and reduces the remaining capacity $B$ by at least $\delta$. Inserting any further items (of either types) results in a negative expected profit:
$\operatorname{\mathbb{E}}[\text{inserting item of type $i$}\;|\;B\leq 1-\delta]$ $\leq\Pr[s_{i}\neq 1]\cdot v_{i}-c_{i}$
@@
We conclude the proof by considering an input instance $I$ with $n>\frac{1}{\epsilon}$ items of type 1 and $n^{2}-n$ items of type 2. The optimal policy does as well as the following policy: it will attempt to insert items of type 1, and stop if an item’s size instantiates $s_{i}\neq 0$, or if no more items of type 1 remain. The expected number of items successfully inserted by OPT is $\frac{1}{2\epsilon}$, each item contributing $\epsilon$ net profit, resulting in
@@
Since $\mathcal{A}$’s only strategy is to randomly select an item to insert, with probability $1-\frac{1}{n}$ it begins by inserting a “bad” item. By taking $n\to\infty$, the expected profit of $\mathcal{A}$ approaches $\epsilon$:
@@
6 $O(1/\delta)$-approximation for Bounded-Size Jobs
In this section, we assume that all jobs are size-bounded with high probability; specifically, we assume that for some $\delta\in(0,1)$, for all $i,j$,
$\Pr_{s_{i}^{j}\sim F_{i}^{j}}[s_{i}^{j}>1-\delta]\leq\frac{1}{2n\alpha}.$
By Lemma 2.1, we focus on providing an approximately optimal policy for the MSKC problem.
As a first step, we analyze a modified problem, where there is no bound on the job sizes, and if a job exceeds the deadline, the principal collects the job’s reward, and then the process terminates. For this case, we give a policy that approximates the profit of the optimal policy that does not receive the reward of the job that overflows. We then show how to transform this policy into a policy that handles bounded-size jobs and does not collect the reward of overflowing jobs.
Our algorithm for the case where jobs can overflow, SKC-OF, runs SKC with budget $\delta$ and a different threshold $\mathcal{T}=\Phi(\delta)/4$. Note that, unlike the SKC algorithm, this threshold is independent of $\alpha$. We denote by $\textbf{SKC-OF}(t)$, the algorithm that runs SKC with a budget $t$, and a threshold $\mathcal{T}=\Phi(t)/4$. Specifically, $\textbf{SKC-OF}=\textbf{SKC-OF}(\delta)$. We have the following.
Lemma 6.1.
For any constant $c\geq 1$,
@@ 6 $O(1/\delta)$-approximation for Bounded-Size Jobs
In case $w_{\max}\geq\Phi(t)/4$, we only take $w_{\max}$ and obtain a profit of
where the last inequality follows Lemma 3.7. Therefore, assume $w_{\max}<\Phi(t)/4$.
Recall that $\textbf{SKC-OF}_{k}$ is the profit of item $k$ in the SKC-OF algorithm, as defined in Section 3. In our case, the profit from item $k$ will be as large as in the MSKC instance, and will not incur a cost in case of an overflow. Therefore, by adapting Eq. (3), our bound on $\textbf{SKC-OF}_{k}$ will be
We adapt the bound on SKC-OF from Eq. (4) accordingly to be
We now turn to the bounded job size case. Our algorithm for this case, SKC-BOUND, will run SKC-OF, and pretend that the budget is $\delta$, i.e. it will terminate if the total size of jobs exceeds $\delta$.
Theorem 6.2.
Proof.
If all items are instantiated with sizes $s_{i}^{j}<1-\delta$, then since the policy terminates when we exceed the budget of $\delta$, the policy never overflows, achieving the expected value of SKC-OF. However, if $s_{i}^{j}>1-\delta$ for some $i,j$, the policy terminates and an overflow may occur, resulting in a loss of at most $\alpha w_{\max}$. We have that
$\geq\Phi(\delta)/4-\alpha w_{\max}\cdot n\cdot\frac{1}{2n\alpha}$
$\geq\Phi(\delta)/4-\Phi(\delta)/8$
$=\Phi(\delta)/8.$
The theorem follows from Lemma 3.7,
We note that SKC-BOUND determines a non-adaptive order of jobs-effort levels (as it applies SKC), but it applies an adaptive stopping rule, as it stops after the total realized sizes first exceed $\delta$. In Appendix B, we show that even for bounded-sized jobs, any non-adaptive policy is bounded away by a factor $\Omega(\sqrt{\alpha})$ from the optimal policy (Theorem B.1).
In Appendix C, we show that if indeed the principal gets the reward of any job that is initiated before the deadline (even if it finishes after the deadline), then we have a constant adaptivity gap Lemma C.1.
7 Conclusion and Future Directions
In this work, we initiate the study of contract design where rewards are fixed, but the principal incentivizes the agents to invest effort in finishing their jobs on time. We identify an economically meaningful Inverse of the Return-on-Investment parameter, $\alpha$, which captures many aspects of the problem’s complexity. Dean et al. [11] show a constant adaptivity gap, while we show that in our case, the adaptivity gap is $\Theta(\alpha)$. As the optimal adaptive contract is intractable, a remaining open problem is whether one can efficiently compute an adaptive contract whose profit is a good approximation to the optimal adaptive contract.
In Section 4.2, we consider a policy with limited adaptivity: the order is non-adaptive, but the policy stops processing jobs once it expects a negative profit. We show that for such policies, their profit can be a factor $\Omega(\alpha^{1/3})$ away from the optimal profit. As our algorithm in Section 3 guarantees an $O(\alpha)$-approximation, this leaves a gap between the lower and upper bound for such policies. It is also interesting to consider other limited-adaptivity policies, e.g., a policy that predetermines the agents’ order, but not their effort levels. Another direction is quantifying the profit loss from using a family of restricted contracts, such as contracts that do not pay agents who finish after the deadline.
Finally, many additional stochastic optimization problems, such as stochastic routing and others, have natural strategic variants were the uncertainty is controlled by an agent’s actions. E.g., in routing, service times are likely to result from a strategic choice of an economic entity. Marrying the classic algorithms developed for such problems with contractual incentives is an intriguing avenue for future research.
Appendix A LP Overconfidence: An $\Omega(\alpha)$ Overestimation of the True Profit
In this section, we provide an instance of the MSKC problem where the value of the LP introduced in Eq. (2) is larger by a factor $\Omega(\alpha)$ from the obtainable value of any policy. This is in stark contrast to the case without costs, where Dean et al. [11] show a non-adaptive policy who’s reward is a constant factor away from the value of the analogous LP. As our non-adaptive policy in Section 3 gives $O(\alpha)$-approximation to the value of the LP in Eq. (2), this approximation is the best possible with respect to the value of the LP.
In our instance, we have an infinite number of identical jobs with value $1$ and cost $1-\epsilon$. The size-distribution of the jobs is
By Definition 2.6, the value of the IOR parameter is $\alpha=\frac{1-\epsilon}{\epsilon}=O(1/\epsilon)$. For every job $i$, we have $\mu_{i}=(1-\epsilon)\cdot\epsilon^{n}+\epsilon\cdot 1=\epsilon^{n}+\epsilon-\epsilon^{n+1}\leq 2\epsilon$ and $w_{i}=\epsilon$. Setting $x_{i}=$ for $1/2\epsilon$ is a feasible assignment to the LP, which implies that $\Phi(1)\geq\frac{1}{2\epsilon}\cdot\epsilon=1/2.$ We next bound the reward of any policy for this instance.
For any policy, after executing the first job, we have that the any job fits with probability at most $1-\epsilon$; thus, we have that the benefit from trying to execute any job after the first one is at most $(1-\epsilon)\epsilon-\epsilon\cdot(1-\epsilon)=0.$ Therefore, the policy that only executes one job is optimal, and it obtains a reward of $\epsilon\leq\frac{\epsilon}{2}\Phi(1)=\Phi(1)/\Omega(\alpha).$
Appendix B The Need for Adaptivity for Bounded-Size Jobs
Theorem B.1.
There exists an instance $I$ of the Knapsack Contracts problem such that
Proof.
Let item:
Property Item
Reward $v$ 22
Cost $c$ $2-\epsilon^{2}$
Size $s$ dist. $s_{1}=\begin{cases}\epsilon^{n},&\text{w.p. }1-\epsilon,\\ \frac{1}{2},&\text{w.p. }\epsilon;\end{cases}$
We consider the case where $S^{\prime}<\frac{1}{2}$, and compute the expected profit from each job.
@@ Appendix B The Need for Adaptivity for Bounded-Size Jobs
$\operatorname{\mathbb{E}}[w|S^{\prime}<\frac{1}{2}]=v\cdot\Pr[s\leq S^{\prime}]-c=2\cdot(1-\epsilon)-(2-\epsilon^{2})=2-2\epsilon-2+\epsilon^{2}=-2\epsilon+\epsilon^{2}<0$
In expectation, we can include approximately $1/\epsilon$ jobs before reaching a state where $S^{\prime}<\frac{1}{2}$ so, the expected number of jobs executed by the optimal stopping-time adaptive policy is $1/\epsilon$, and therefore, it obtains an expected profit of $\frac{1}{\epsilon}\cdot\epsilon^{2}=\epsilon.$
Consider now a non-adaptive policy. This policy decides on the number of jobs to execute, and ends whether it overflows or completes the number of jobs it sets to execute. Consider first a policy that only executes one job. Its reward is exactly $v-c=\epsilon^{2}$. Let $\mathrm{profit}(i)$ be the expected profit from the $i$th job in the sequence. Since the first job always finishes before the deadline, it is clear that $\mathrm{profit}(1,2)=c-v=\epsilon^{2}$. Consider now $\mathrm{profit}(i)$ for $i>2$. We have that
$\mathrm{profit}(i)$ $=$ $\operatorname{\mathbb{E}}[\mathrm{profit}(i)|s_{j}=\epsilon^{n}\ \forall j<i]\cdot\Pr[s_{j}=\epsilon^{n}\quad\forall j<i]$
$+\operatorname{\mathbb{E}}[\mathrm{profit}(i)|\exists!j<i:\ s_{j}>\epsilon^{n}]\cdot\Pr[\exists!j<i:\ s_{j}>\epsilon^{n}]$
$=$ $(v-c)\cdot(1-\epsilon)^{i-1}+((v-c)\cdot(1-\epsilon)-c\cdot\epsilon)\cdot\Pr[\exists!j<i:\ s_{j}>0]$
$=$ $\epsilon^{2}\cdot(1-\epsilon)^{i-1}+((1-\epsilon)v-c)\cdot\Pr[\exists!j<i:\ s_{j}>0]$
$=$ $\epsilon^{2}\cdot(1-\epsilon)^{i-1}+(\epsilon^{2}-2\epsilon)\cdot(1-\epsilon)^{i-2}\cdot\epsilon$
$\leq$ $\epsilon\cdot(1-\epsilon)^{i-2}(\epsilon^{2}-\epsilon)$
$<$ $0.$
In the above, the first inequality follows from the fact that $\epsilon^{2}-2\epsilon$ is negative, and that the probability that exactly one of the first i1i-1 jobs has size greater than $\epsilon^{n}$ is larger than the probability that only job i1i-1 has size greater than $\epsilon^{n}$. By Definition 2.6, the value of the IOR parameter is $\alpha=\frac{2-\epsilon^{2}}{\epsilon^{2}}=O(1/\epsilon^{2})$.
Therefore, we have that a non-adaptive policy only loses from having a sequence larger than 2, and the optimal non-adaptive policy has a profit of $2\epsilon^{2}$. We conclude that for this instance, we have $\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}=\frac{1}{2\epsilon}=O(\sqrt{\alpha}).$
Appendix C Constant Approximation for MSKC with Overflows
We first provide an upper-bound for expected profit of the optimal adaptive policy that can overflow as well. Recall the definition of $\Phi$ given in Eq (2). Let $\mathrm{ADAPT}_{OF}$ be the optimal adaptive policy that can overflow.
Lemma C.1.
For every instance of the knapsack contract problem $I$, $\mathrm{ADAPT}_{OF}(1)\leq 2\Phi(1)+w_{max}$.
Proof.
let $\mathcal{S}$ be the random set of item choices that the policy inserts, and $w_{ij}=v_{i}\cdot\Pr_{{s_{i}^{j}\sim F_{i}^{j}}}[s_{i}^{j}\leq 1]-p_{i}^{j}$ (like in Section 3), $w_{\max}=\max_{(i,j)\in\mathcal{C}}w_{ij}$ (Algorithm 1). $\Phi(t)$ is defined in Eq. (2), .
$\leq$ $\sum_{ij\in\mathcal{S}}w_{ij}\cdot x_{ij}+w_{max}$
$\leq$ $\Phi(2)+w_{max}$
$\leq$ $2\Phi(1)+w_{max}.$
Here the first inequity follows since the probability $s_{i}>1$ is independent from the fact $(i,j)\in\mathcal{S}$. The last inequality follows from the fact that, at each run of the adaptive policy, at most one element has a size larger than 1, and therefore the sum of probabilities of the right summand is at most 1. By Lemma 3.6, we know that $\operatorname{\mathbb{E}}[\mu(\mathcal{S})]\leq 2$, therefore, we get $\sum_{ij\in A}w_{ij}\cdot x_{ij}\leq\Phi(2)$. ∎
Theorem C.2.
$\textbf{SKC}_{OF}$ is a 9-approximation algorithm to $\mathrm{ADAPT}_{OF}$.
Proof.
By Equations (7),(8):
$\frac{ADAPT_{OF}}{\textbf{SKC-OF}}$ $\leq$ $\frac{2\Phi(1)+w_{max}}{\Phi(1)/4}.$
If $w_{max}>\Phi(1)/4$: so $\textbf{SKC-OF}=w_{max}$ and:
$\frac{ADAPT_{OF}}{\textbf{SKC-OF}}$ $\leq$ $\frac{8w_{max}+w_{max}}{w_{max}}=9.$
Else:
$\frac{ADAPT_{OF}}{\textbf{SKC-OF}}$ $\leq$ $\frac{2\Phi(1)+\Phi(1)/4}{\Phi(1)/4}=9.$
Appendix D Ordered Adaptive Contracts
In many real-world scenarios, the principal lacks the flexibility to reorder agents at will. For example, this can be due to procedural requirements, like a fixed sequence of suppliers in a supply chain, or pre-determined project plan with sequential dependencies. In these settings, the principal’s adaptive power is limited to deciding, after observing the outcome of each task, whether or not to continue with the next agent, and if so – which effort level to incentivize. We define the Ordered Adaptive Contracts model:
Definition D.1 (Ordered Adaptive Contracts Model).
In the Ordered Adaptive Contracts model, the principal is given a fixed ordering $\sigma=(1,2,\ldots,n)$ of the $n$ agents. For each agent $i$ in order, the principal offers a contract $t_{i}$, and the agent chooses an effort level $j$ to maximize their utility. The agents execute their jobs sequentially in the given order. The principal observes the completion time of each job and may adaptively decide whether to proceed to the next agent or stop, but cannot change the order. The objective is to maximize the expected net profit (total value from completed jobs before the budget is exhausted minus total payments to agents), given the fixed order $\sigma$.
Since the order is fixed, we are interested in comparing our policy’s performance to the performance of the optimal policy operating on the same order of agents. By using the same reduction to the multi-agent contract problem as shown in Lemma 2.1, we focus on solving the matching stochastic knapsack with costs problem. In a similar manner to the algorithm that finds an ordered adaptive policy presented in [11] and to the PTAS presented in [5] for the regular Stochastic Knapsack variant (single-choice, no-cost), we use a resource augmentation approach: the policy is allowed a knapsack of size $(1+\epsilon)$, while the optimal adaptive policy has a knapsack of size $1$. Our algorithm runs in time $O(\frac{n^{3}m}{\epsilon^{2}})$, offering a tradeoff between the augmented size of the knapsack and the running time of the algorithm.
We use a dynamic programming algorithm with discretized sizes. We denote by $\mathcal{D}(i,k)$ the optimal expected value that can be obtained from inserting items i,,ni,...,n, with $\delta k$ capacity left in the knapsack, for some small $\delta>0$ to be determined later. We are interested in computing $\mathcal{D}(1,\lfloor\frac{1}{\delta}\rfloor)$. If all of the real sizes are indeed multiples of $\delta k$, then one can obtain an optimal policy by computing:
$\mathcal{D}(i,k)=\max\left\{\mathcal{D}(i+1,k),\max_{j\in[m]}\left\{v_{i}\cdot\Pr[s_{i}^{j}\leq\delta k]-p_{i}^{j}+\sum_{t=0}^{k}{\mathcal{D}(i+1,k-t)\cdot\Pr[s_{i}^{j}=t\delta]}\right\}\right\}.$
With the base cases of $\mathcal{D}(n+1,k)=0$ for all $k$, and $\mathcal{D}(i,k)=0$ for $k<0$.
However, when the distributions are arbitrary, a slight deviation from the discretization can result in an arbitrarily bad solution. To overcome this, we must utilize the fact our policy is allowed a knapsack of size $1+\epsilon$. We do so by rounding down the true distribution at each step to the nearest multiple of $\delta$. By picking a small enough $\delta$, we ensure that the overall rounding error is not too large.
Theorem D.1.
There exists an $O(\frac{n^{3}m}{\epsilon^{2}})$ time algorithm that finds a policy for the ordered adaptive knapsack with costs problem, where the policy has a knapsack of size $(1+\epsilon)$, and the optimal adaptive policy has a knapsack of size $1$. The policy achieves an expected profit of at least the profit of the optimal policy with a knapsack of size $1$.
Proof.
Let OAC be our algorithm and $\mathrm{ADAPT}$ be the optimal adaptive policy. We set the discretization parameter $\delta=\epsilon/n$.
First, consider a modified problem where all item sizes $s_{i}^{j}$ are rounded down to the nearest multiple of $\delta$: $\hat{s}_{i}^{j}=\delta\lfloor s_{i}^{j}/\delta\rfloor$. Let $\widehat{\mathrm{ADAPT}}$ be the optimal expected profit for this problem with rounded-down sizes. Since for any policy, the probability of successfully fitting a set of items can only increase when item sizes are reduced, the optimal value for the rounded-down problem is an upper bound on the optimal value for the original problem. Thus, $\widehat{\mathrm{ADAPT}}\geq\mathrm{ADAPT}.$
Our algorithm solves the dynamic program $\mathcal{D}$ described above for the rounded-down sizes. This DP computes the exact optimal policy for the rounded-down problem. Denote by $\varphi_{\textbf{OAC}}$ the policy derived from this DP. When given a knapsack of capacity $1$, this policy achieves the optimal value for the rounded-down problem. Let $\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)$ be a random variable of the profit of the policy with the original capacity and rounded-down job sizes. We have that $\operatorname{\mathbb{E}}[\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)]=\widehat{\mathrm{ADAPT}}$. Combining these, we have $\operatorname{\mathbb{E}}[\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)]\geq\mathrm{ADAPT}.$
Now, we analyze the performance of our algorithm, which executes the policy $\varphi_{\textbf{OAC}}$ on the original item sizes $s_{i}^{j}$ with an augmented capacity of $1+\epsilon$. Consider any execution path of $\varphi_{\textbf{OAC}}$ where a sequence of items and choices $\mathcal{S}\subset[n]\times[m]$ is attempted and successfully completed. According to the policy $\varphi_{\textbf{OAC}}$ (which was computed for capacity 1 and sizes $\hat{s}$), the sum of the rounded-down sizes must be no more than 1:
$\sum_{(i,j)\in\mathcal{S}}\hat{s}_{i}^{j}\leq 1.$
The true total size of these items is $\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}$. We know that for each item, $s_{i}^{j}=\hat{s}_{i}^{j}+\eta_{i}^{j}$, where $0\leq\eta_{i}^{j}<\delta$ is the rounding error. The total size in the original problem is:
$\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}=\sum_{(i,j)\in\mathcal{S}}\hat{s}_{i}^{j}+\sum_{(i,j)\in\mathcal{S}}\eta_{i}^{j}.$
Since at most $n$ items can be placed in the knapsack, the total accumulated rounding error is bounded:
$\sum_{(i,j)\in\mathcal{S}}\eta_{i}^{j}\leq n\cdot\delta=n\cdot(\epsilon/n)=\epsilon.$
Therefore, the true total size of the items is:
$\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}\leq\sum_{(i,j)\in\mathcal{S}}\hat{s}_{i}^{j}+\epsilon\leq 1+\epsilon.$
This shows that any sequence of items that fits within capacity 1 in the rounded-down instance is guaranteed to fit within the augmented capacity of $1+\epsilon$ in the original instance. Since the policy $\varphi_{\textbf{OAC}}$ makes decisions based on the sequence of outcomes (which are independent of the rounding), every successful path under capacity 1 with sizes $\hat{s}$ is also a successful path under capacity $1+\epsilon$ with sizes $s$. Moreover, it crucially follows that an item can cause an overflow under capacity $1+\epsilon$ with sizes $s$ only if it causes an overflow under capacity 1 with sizes $\hat{s}$. Therefore, the expected profit can only be higher:
$\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi_{\textbf{OAC}},\text{capacity}=1+\epsilon)]\geq\operatorname{\mathbb{E}}[\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)].$
Chaining the inequalities, we get the final result:
$\textbf{OAC}=\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi_{\textbf{OAC}},\text{capacity}=1+\epsilon)]\geq\operatorname{\mathbb{E}}[\widehat{\mathrm{profit}}(\varphi_{\textbf{OAC}},\text{capacity}=1)]\geq\mathrm{ADAPT}.$
The dynamic program $\mathcal{D}$ has $O(n)\cdot O(n/\epsilon)$ states, and computing each state takes up to $O(m\cdot n/\epsilon)$ time. This results in a running time of $O(\frac{n^{3}m}{\epsilon^{2}})$.
Appendix E Omitted Proofs from Section 3

### v2

@@ Appendix C $\Omega(\alpha)$ Informational Gap
In this section we establish a lower bound that exposes a fundamental information gap in the problem of Stochastic Knapsack with Costs. Previous constant-factor approximation algorithms for Stochastic Knapsack (without costs) have used limited distributional information, specifically the expected (or truncated) sizes and the probability of fitting within the current capacity. In Proposition C.1 we demonstrate that this information is insufficient. We show that any algorithm, even one that finds a fully-adaptive policy, which relies solely on such limited data, cannot achieve a constant-factor approximation in the presence of costs. More generally, we prove in Theorem C.2 that even providing the first $k$ moments of each distribution for any constant $k$ is insufficient.
C.1 Warm-Up: First Moment
Proposition C.1.
@@ C.1 Warm-Up: First Moment
Note that in this instance the expected values and the truncated expected values are identical: $\operatorname{\mathbb{E}}[s_{1}]=\operatorname{\mathbb{E}}[s_{2}]=\mu_{1}=\mu_{2}=\epsilon$ and $q_{1}(1)=q_{2}(1)=1$, hence, algorithm $\mathcal{A}$ cannot distinguish between the two items. Inserting one item of type 2 results in a net utility of $\epsilon$, and reduces the remaining capacity $B$ by at least $\epsilon/2$. Inserting any further items (of either types) results in a negative expected utility:
@@
Now, consider an input instance $I$ with $n>\frac{1}{\epsilon}$ items of type 1 and $n^{2}-n$ items of type 2. The optimal policy will attempt to insert items of type 1, and stop if an item’s size instantiates $s_{i}=1$, or if no more items of type 1 remain. The expected number of items successfully inserted by OPT is $\frac{1}{\epsilon}$, each item contributing $\epsilon$ net utility, resulting in
@@
Since $\mathcal{A}$’s only strategy is to randomly select an item to insert, with probability $1-\frac{1}{n}$ it begins by inserting a “bad” item. By taking $n\to\infty$, the expected utility of $\mathcal{A}$ approaches $\epsilon$:
@@ C.2 $k$ Moments
Inserting a single item of type 2 results in a net utility of $\epsilon$, and reduces the remaining capacity $B$ by at least $\delta$. Inserting any further items (of either types) results in a negative expected utility:
$\operatorname{\mathbb{E}}[\text{inserting item of type $i$}\;|\;B\leq 1-\delta]$ $\leq\Pr[s_{i}\neq 1]\cdot v_{i}-p_{i}$
@@
We conclude the proof by considering an input instance $I$ with $n>\frac{1}{\epsilon}$ items of type 1 and $n^{2}-n$ items of type 2. The optimal policy does as well as the following policy: it will attempt to insert items of type 1, and stop if an item’s size instantiates $s_{i}\neq 0$, or if no more items of type 1 remain. The expected number of items successfully inserted by OPT is $\frac{1}{2\epsilon}$, each item contributing $\epsilon$ net utility, resulting in
@@
Since $\mathcal{A}$’s only strategy is to randomly select an item to insert, with probability $1-\frac{1}{n}$ it begins by inserting a “bad” item. By taking $n\to\infty$, the expected utility of $\mathcal{A}$ approaches $\epsilon$:
@@
Appendix D Omitted Proofs from Section 3
D.1 Proof of theorem 3.1
@@ Appendix D Omitted Proofs from Section 3
For a constant $\alpha$, the claim is trivially true. We prove the claim for every $\alpha>3$. We build an instance $I$ as follows. Consider some $\epsilon<\frac{1}{2}$ and any $\gamma\in(0,\tfrac{1}{2})$, where $\epsilon$ will be used to determine $\alpha$. Our instance has $n$ items for $n=\lceil\frac{1}{\epsilon}\rceil$. Each item has value $v=2$ and cost $p=2-\epsilon$.
The size distribution of item $i\in[n]$ is
$s_{i}\;=\;\begin{cases}a_{i},&\text{with probability }1-\epsilon,\\[4.0pt] b_{i},&\text{with probability }\epsilon,\end{cases}$ where $\qquad a_{i}=\gamma^{\,n-i+1},\qquad b_{i}=1-\displaystyle\sum_{j=1}^{i-1}a_{j}$
By Definition 2.1, we have $\alpha=\frac{2-\epsilon}{\epsilon}\;=\;\Theta\!\left(\frac{1}{\epsilon}\right).$ Note that for every $\alpha>3$, we can choose an $\epsilon<1/2$ in order to set the IOR of our instance to be $\alpha$.
Since for every $i$, $\sum_{j=1}^{i-1}a_{j}+b_{i}=1,$ a stopping-adaptive policy can safely insert the items sequentially until it encounters the first item $i$ whose realized size equals $b_{i}$, and then halt.
We show that since $b_{i}+a_{i^{\prime}}>1$ for every pair $i<i^{\prime}$, attempting to insert item $i^{\prime}$ without considering the realization of earlier items’ size yields negative expected utility. Consequently, the optimal $\mathrm{NON\_ADAPT}$ policy executes at most a single item. Our proof combines these two facts to yield the desired bound.
Formally, consider a stopping-adaptive policy that inserts items in increasing index order ($i=1,2,\ldots$) until it encounters the first item whose size realizes as $s_{i}=b_{i}$, and then stops. Let $X$ denote the number of items the policy inserts. We have
Since $n\geq 1/\epsilon$, we get that $\operatorname{\mathbb{E}}[X]\geq\frac{1-e^{-1}}{\epsilon}$. Therefore the expected utility is at least
@@
$\mathrm{STOP\_ADAPT}(I)\;=\;\operatorname{\mathbb{E}}[X]\,(v-p)\;\geq\;\frac{1-e^{-1}}{\epsilon}\,\cdot\,\epsilon\;=\;1-e^{-1}.$
Consider now a non‑adaptive policy. This policy fixes the number of items to insert, $k$, and an ordering $\pi=i_{1},i_{2},\ldots,i_{k}$. For each $\ell\in[k]$, let $\mathrm{utility}(\ell)$ denote the utility from the $\ell$‑th item in the sequence. Clearly, $\operatorname{\mathbb{E}}[\mathrm{utility}(1)]=\epsilon$ for every non‑empty sequence of items. We will show that for every sequence with $k>1$ we have $\operatorname{\mathbb{E}}[\mathrm{utility}(k)]<0$, which implies that it is always better to execute at most a single item, since we can iteratively remove items and only improve the expected utility of the policy.
Given a sequence $\pi=i_{1},i_{2},\ldots,i_{k}$ of items, consider first the case in which $i_{\ell}>i_{k}$ for some earlier item in the sequence ($\ell<k$). If $i_{k}$ is reached and its size realises as $s_{i_{k}}=b_{i_{k}}$, the item overflows, because $s_{i_{\ell}}+s_{i_{k}}\geq a_{i_{\ell}}+b_{i_{k}}>1$. Hence
[…]
$\leq\Pr[i_{k}\text{ is reached}]\Bigl(\Pr[s_{i_{k}}=b_{i_{k}}](-p)+\Pr[s_{i_{k}}=a_{i_{k}}](v-p)\Bigr)$
$=\Pr[i_{k}\text{ is reached}]\bigl(\epsilon(-p)+(1-\epsilon)(v-p)\bigr)$
$=-\,\epsilon\,\Pr[i_{k}\text{ is reached}]\;<\;0.$
Now consider the case where $i_{\ell}<i_{k}$ for all $\ell<k$. Let $j=\arg\max_{\ell<k}i_{\ell}$, the largest‑index item among the first k1k-1 inserted. Note that, $i_{j}<i_{k}$ by our assumption. We have that
$\operatorname{\mathbb{E}}[\mathrm{utility}(k)]$ $=$ $\operatorname{\mathbb{E}}[\mathrm{utility}(k)|i_{k}\mbox{ is reached}]\cdot\Pr[i_{k}\mbox{ is reached}]$
$=$ $\operatorname{\mathbb{E}}[\mathrm{utility}(k)|s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k]\cdot\Pr[s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k]$
$+\operatorname{\mathbb{E}}[\mathrm{utility}(k)|i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
$\ \ \cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
$=$ $(v-p)\cdot(1-\epsilon)^{k-1}+(-p)\cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
$=$ $\epsilon\cdot(1-\epsilon)^{k-1}+(-p)\cdot\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]$
$\leq$ $\epsilon\cdot(1-\epsilon)^{k-1}+(-p)\cdot\Pr[s_{i_{j}}=b_{i_{j}}\ \wedge\ (s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k,\ell\neq j)]$
$=$ $\epsilon\cdot(1-\epsilon)^{k-2}(1-\epsilon-2+\epsilon)$
$=$ $-\epsilon\cdot(1-\epsilon)^{k-2}<0.$
In the above, the third equality follows since $i_{k}$ always overflows in case some $s_{i_{\ell}}=b_{i_{\ell}}$ for some $i_{\ell}<i_{k}$, as $b_{i_{\ell}}+a_{i_{k}}>1$. The first inequality follows from the fact that p-p is negative, and $\Pr[i_{k}\mbox{ is reached}\ \wedge\ \exists\ell<k:\ s_{i_{\ell}}=b_{i_{\ell}}]\geq\Pr[s_{i_{j}}=b_{i_{j}}\ \wedge\ (s_{i_{\ell}}=a_{i_{\ell}}\quad\forall\ell<k,\ell\neq j)]$ since for the event $s_{i_{j}}=b_{i_{j}}$ and for all other $\ell<k$, $s_{i_{\ell}}=a_{i_{\ell}},$ we have that $i_{k}$ is reached. Since $\sum_{\ell<k,\ell\neq j}a_{i_{\ell}}\leq\sum_{\ell=1}^{i_{j}-1}a_{\ell}=1-b_{i_{j}},$ there will be no overflowing item before $i_{k}$ is executed.
We conclude that the optimal non-adaptive policy only inserts a single item, and obtains a utility of $\epsilon$, while the optimal stopping-adaptive policy obtains a utility of 1, as argued above. Therefore, $\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}\geq\frac{1-e^{-1}}{\epsilon}=\Omega(\alpha).$
D.2 Proof of lemma 3.4